Draw l, through (-4, 0) and (0, 6). What is the slope of any line parallel to l1? Draw l2 through the origin and parallel to l1.
Answer by user_dude2008(1862)

m=(y2-y1)/(x2-x1)=(6-0)/(0--4)=6/4=3/2

slope of l1 and l2 is 3/2